Terrence and Quentin picked apples. Terrence picked 3 times as many pounds of apples as Quentin. Together they picked 28 pounds of apples. How many pounds of apples did Terrence pick?
Answer by richwmiller(17219)

t+q=28
t=3*q
3*q+q=28
4*q=28
q=7
t=3*7
t=21
check
21+7=28
ok